OpenStudy (anonymous):

If a sphere has a radius of 2 untis ,then what is the volume of half of the sphere? A.) 32/3π cubic units B.)16π cubic units C.)16/3π cubic units D.)64/3π cubic units

OpenStudy (anonymous):

Volume of Hemisphere=(4/6) Pi r^3 Thus, (4/6)*Pi*2^3=(16/3)*Pi

OpenStudy (anonymous):

i explained this to you already. use the formula i provided earlier

OpenStudy (anonymous):

i though this was a different kind of problem

OpenStudy (anonymous):

are you sure i use the same formula

OpenStudy (anonymous):

yes, volume of the whole sphere is 4/3 pi r ^3. so volume of half the sphere is half of 4/3pi r^3

OpenStudy (anonymous):

man, you just need to relax, everything is better than you could possibly imagine :) so, just go outside, chill a bit, ride a bike or something. You could also go horseriding. So mate, nothing to worry about, forget your homework. School is just there to confuse you. :)

OpenStudy (anonymous):

okay i thought i was doing a different problem so i needed a different formula...thanks again

OpenStudy (anonymous):

just pause for a second and think about what you are trying to solve. and you are welcome BTW. wayofaway's answer is correct.

OpenStudy (anonymous):

yes thanks and i didn't even see a answer lol

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

hey can you check one of my answer on another problem?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

(-3+16i)+(-6-4i) A.)-9+12i B.)9+12i C.)3+200i D.)-9+20i I think its A but can someone check me?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

yup. add the real parts and add the imaginary parts separately.

OpenStudy (anonymous):

yayyy haha thats what my teacher said lol

OpenStudy (anonymous):

okay great!